You observe a nominal rate of 14.23% at the same time you learn that the real rate is 2.12%. What is the inflation rate?
Answer #1

The Fisher equation, which shows the exact relationship between nominal interest rates, real interest rates, and inflation, is:

(1 + R) = (1 + r)(1 + h)

(1 + 0.1423) = (1 + 0.0212)(1 + h)

h = [(1 + 0.1423) / (1 + 0.0212)] – 1

h = 0.1186 or 11.86%